# Hound: Locating Cryptographic Primitives in Desynchronized Side-Channel Traces Using Deep-Learning

### v1

@@ IV-A Experimental Setup: Hardware and Software
We utilized the NewAE CW305 development board [14] as our validation platform, designed specifically to facilitate high-precision, low-noise side-channel measurements. This board incorporates a Xilinx Artix7-100 FPGA. Power traces were captured using a Picoscope 5244d digital sampling oscilloscope (DSO), operating at a sampling rate of 125 Msamples/s and a resolution of 12 bits. For our reference computing platform, we employed a 32-bit RISC-V System-on-Chip implemented on the FPGA. Taking steps from [23], we implemented the Dynamic Frequency Scaling mechanism coupled with a true random number generator (TRNG) to deliver a novel actuator that can randomize the operating frequency thus introducing non-negligible deformations in the collected side-channel traces. The TRNG generates a random number used to select a frequency from a pool of 760 available frequencies ranging from 5 MHz to 100 MHz, with a step of 125 kHz. Notably, DFS remained enabled throughout the entire experimental phase. Consequently, a new frequency is requested as soon as a configuration is locked. As a result, each cryptographic primitive experiences a variable number of clock frequency reconfigurations, e.g., an average of 41 reconfigurations per AES cryptosystem, leading to significant trace deformation. As the cryptographic primitives of choice, we selected the constant-time, unprotected version of three ciphers, i.e., AES-128, Clefia-128, and Camellia-128, from the OpenSSL software codebase [27], and a masked version of Tiny-AES-128 [28].
@@ IV-B Training Evaluation
Each neural network underwent training for $25$ epochs, where an epoch encompasses a complete iteration over the training set. The Adam optimizer [29] was employed to minimize the cross-entropy loss function. The validation error was assessed after each epoch, and the network with the lowest error was chosen for further evaluation. Subsequently, this network was employed to assess the performance of unseen traces during the inference phase. Since DFS introduces significant trace deformation with high variability, we noticed how CNNs that did not reach an accuracy of at least 99% were not always able to classify the CPs in the sliding windows procedure correctly. Hyperparameters such as training batch size, learning rate, and dropout were carefully chosen to address this issue.
@@ IV-C Inference Evaluation
$IoU_{f}=\frac{|P_{i}\cap GT_{i}|}{|P_{i}\cup GT_{i}|}$ (1)
Since Hound focuses on pinpointing the CP’s beginning within the trace, the IoU metric disregards the CP’s ending point. The ground truth ending point is always assumed to be accurate. Table I shows the mean and standard deviation of IoU, where a higher score corresponds to a better segmentation. While IoU calculation requires a label that precisely indicates the CP’s beginning, it is essential to note that it is never used during the inference pipeline itself. It is solely used for calculating these metrics.
We conducted tests on the inference pipeline for each cipher using consecutive cipher executions and encryptions interleaved with random applications. Evaluating a sequence of consecutive CP executions assesses the robustness of the proposed method in locating CPs when they occur sequentially. Conversely, executing CPs mixed with noisy applications evaluates the effectiveness of the proposed method in locating CPs within a heterogeneous application scenario. The segmentation hits score achieved 100% for every cryptographic algorithm in both scenarios, i.e., consecutive encryption and interleaved with noisy applications, consistently managing to identify all executions. The IoU score, on the other hand, ranges from 91.92% to 98.46% for the different ciphers. The results demonstrate the robustness of the proposed methodology in locating CPs with high accuracy and precision, even in the presence of DFS-induced trace deformation.
@@
Figure 5: Example of Hound inference pipeline applied to a side-channel trace affected by DFS that contains 5 AES executions mixed with noise applications.
TABLE II: Segmentation and SCA results targeting AES-128. Reported results considers the presence (or not) of noise applications interleaved with the CPs in the side-channel trace.
Hiding Noise Hits SCA
@@ IV-D The Complete Attack Flow
Figure 5 shows the results of applying each step of the proposed inference pipeline to locate the CPs into a portion of a trace containing five executions of AES-128, i.e., the CP. The entire side-channel trace contains 367 CPs collected from a platform implementing a randomized dynamic frequency scaling mechanism. Starting from the portion of the side-channel trace in input, Figure 5b depicts the segmentation output of the Sliding Window Classification stage that correctly highlights the presence of five CPs interleaved with noise applications. Notably, it is easy to detect also the end of the AES encryptions by looking at Class 1, as the classifier is trained to recognize the spare part of the CPs. The number of samples in the x-axis of the plot is reduced by a factor of $62$, equal to the value of stride $s$ during the classification. The Screening stage cleans the segmentation output by removing spurious activations of the various classes (see Figure 5c). The time instants locating the beginning of the five CPs are applied to the original side-channel trace, thus allowing an easy time realignement (see Figure 5d).
@@
Comparison against state of the art - Building on the prior evaluation, Table II further stresses the effectiveness of the proposed methodology, detailing the attack results under various scenarios. The target CP is AES-128, we compare our approach against three state-of-the-art proposals, i.e., [17], [18], and [19]. For each state-of-the-art approach, we report the hiding mechanism for which the methodologies were designed by their authors (see the Hiding Mechanism column in Table II). We considered the random execution of the CPs within a set of noisy applications, as well as the continuous execution of the CPs without any noisy application (see the Noise Applications column in Table II). For each analyzed scenario, Hound can correctly identify the beginning of all the CPs in the side-channel trace, leading to a successful SCA attack. In contrast, the three state-of-the-art methodologies fail to locate the CPs in the side-channel trace due to the high DFS obfuscation, and thus, the subsequent side-channel attack is unsuccessful. More specifically, [17] and [18] achieve a 0% hit rate, indicating they cannot locate any CPs in the trace. On the other hand, [19] exhibits anomalously high hit rates of 879% and 1637% for the two scenarios. These excessively high hit-rate values highlights a significant number of false positives determining the subsequent failure of the side-channel attacks.
@@ V Conclusions
This work presented Hound, a novel deep-learning technique for pinpointing cryptographic primitives within highly desyncronized side-channel traces. Unlike existing methods, Hound successfully locates the COs in side-channel traces collected from computing platforms implementing dynamic frequency scaling mechanism as effective hiding countermeasure.
We conducted a comprehensive evaluation to locate both sequences of consecutive cryptographic primitives and those interleaved with other applications. The evaluation considered different cryptographic operations. The experimental results, obtained by running diverse set of applications on an FPGA-based RISC-V processor, validate the effectiveness of our approach in enabling successful side-channel analysis (SCA) attacks. These results also emphasize the shortcomings of current state-of-the-art solutions.

### v3

@@ IV-A Experimental Setup: Hardware and Software
We utilized the NewAE CW305 development board [14] as our validation platform, designed specifically to facilitate high-precision, low-noise side-channel measurements. This board incorporates an AMD Artix7-100 FPGA. Power traces were captured using a Picoscope 5244d digital sampling oscilloscope (DSO), operating at a sampling rate of 125 Msamples/s and a resolution of 12 bits. For our reference computing platform, we employed a 32-bit RISC-V System-on-Chip [26] implemented on the FPGA. Taking steps from [27], we implemented the Dynamic Frequency Scaling mechanism coupled with a true random number generator (TRNG) [28] to deliver a novel actuator that can randomize the operating frequency, thus introducing non-negligible deformations in the collected side-channel traces. The TRNG generates a random number used to select a frequency from a pool of 760 available frequencies ranging from 5 MHz to 100 MHz, with a step of 125 kHz. Notably, DFS remained enabled throughout the entire experimental phase. Consequently, a new frequency is requested as soon as a configuration is locked. As a result, each cryptographic primitive experiences a variable number of clock frequency reconfigurations, e.g., an average of 41 reconfigurations per AES encryption, leading to significant trace deformation. As the cryptographic primitives of choice, we selected the constant-time, unprotected version of three ciphers, i.e., AES-128, Clefia-128, and Camellia-128, from the OpenSSL software codebase [29], and a masked version of Tiny-AES-128 [30].
@@ IV-B Training Evaluation
Each neural network underwent training for $25$ epochs, where an epoch encompasses a complete iteration over the training set. The Adam optimizer [31] was employed to minimize the cross-entropy loss, paired with the one-cycle learning rate scheduler. The validation error was assessed after each epoch, and the network with the lowest error was chosen for further evaluation. Subsequently, this network was employed to assess the performance of unseen traces during the inference phase. Since DFS introduces significant trace deformation with high variability, we noticed how CNNs that did not reach an accuracy of at least 99% were not always able to classify the CPs in the sliding windows procedure correctly. Hyperparameters such as training batch size, learning rate (lr), and dropout were carefully chosen to address this issue (see the CNN hyperparameters column in Table I).
@@ IV-C Inference Evaluation
$IoU_{i}=\frac{|P_{i}\cap GT_{i}|}{|P_{i}\cup GT_{i}|}$ (1)
Since Hound focuses on pinpointing the CP’s beginning within the trace, the IoU metric disregards the CP’s ending point. The predicted ending point is always assumed to be accurate. Table I shows the mean and standard deviation of IoU, where a higher score corresponds to a better segmentation.
We conducted tests on the inference pipeline for each cipher using consecutive cipher executions and encryptions interleaved with random general-purpose applications. The segmentation hits score achieved 100% for every cryptographic algorithm in both scenarios, i.e., consecutive encryption and interleaved with noisy general-purpose applications, consistently managing to identify all executions. The IoU score, on the other hand, ranges from 91.92% to 98.46% for the different ciphers. The results demonstrate the robustness of the proposed methodology in locating CPs with high accuracy and precision, even in the presence of DFS-induced trace deformation.
@@
Figure 5: Example of Hound inference pipeline applied to a side-channel trace affected by DFS that contains 5 AES executions mixed with general-purpose applications.
TABLE II: Segmentation and side-channel attack (SCA) results targeting AES-128. Reported results consider the presence (or not) of general-purpose applications interleaved with the CPs in the side-channel trace.
Hiding General-Purpose Hits SCA
@@ IV-D The Complete Attack Flow
Figure 5 shows the results of applying each step of the proposed inference pipeline to locate the CPs into a portion of a trace containing five executions of AES-128, i.e., the CP. The entire side-channel trace contains 367 CPs collected from a platform implementing a randomized dynamic frequency scaling mechanism. Starting from the portion of the side-channel trace in input, Figure 5b depicts the segmentation output of the Sliding Window Classification stage that correctly highlights the presence of five CPs interleaved with random general-purpose applications. Notably, it is easy to detect also the end of the AES encryptions by looking at Class 1, as the classifier is trained to recognize the spare part of the CPs. The number of samples in the x-axis of the plot is reduced by a factor of $62$, equal to the value of stride $s$ during the classification. The Screening stage cleans the segmentation output by removing spurious activations of the various classes (see Figure 5c). The time instants locating the beginning of the five CPs are applied to the original side-channel trace, thus allowing an easy time realignement (see Figure 5d).
@@
Comparison against state of the art - Building on the prior evaluation, Table II further stresses the effectiveness of the proposed methodology, detailing the attack results under various scenarios. The target CP is AES-128. We compare our approach against three state-of-the-art proposals, i.e., [17], [18], and [19]. For each state-of-the-art approach, we report the hiding mechanism for which the methodologies were designed by their authors (see the Hiding Mechanism column in Table II). We considered the random CPs’ execution within a set of noisy general-purpose applications and their execution without any noisy application (see the General-Purpose Applications column in Table II). For each analyzed scenario, Hound can correctly identify the beginning of all the CPs in the side-channel trace, leading to a successful side-channel attack. In contrast, the three state-of-the-art methodologies fail to locate the CPs in the side-channel trace due to the high DFS obfuscation, and thus, the subsequent side-channel attack is unsuccessful. More specifically, [17] and [18] achieve a 0% hit rate, indicating they cannot locate any CPs in the trace. On the other hand, [19] exhibits anomalously high hit rates of 879% and 1637% for the two scenarios. These excessively high rate values highlight a significant number of false positives determining the subsequent failure of the side-channel attacks.
@@ V Conclusions
We presented Hound, a novel deep-learning technique for pinpointing cryptographic primitives within highly desynchronized side-channel traces. Unlike existing methods, Hound successfully locates the CPs in side-channel traces collected from computing platforms implementing dynamic frequency scaling mechanism as an effective hiding countermeasure.
We conducted a comprehensive evaluation to locate both sequences of consecutive cryptographic primitives and those interleaved with other applications. The evaluation considered different cryptographic primitives. The experimental results, obtained by running a diverse set of applications on an FPGA-based RISC-V processor, validate the effectiveness of our approach in enabling successful side-channel attacks. These results also emphasize the shortcomings of current state-of-the-art solutions.